OpenStudy (anonymous):

Explain the Factor by Grouping method for a trinomial of the form ax²+bx+c where a is a coefficient other than 1

OpenStudy (kenljw):

this is the way I'd do it, maybe not the way taught ax²+bx+c = a(x^2 + (b/a)x + c/a) find the factors and multiply a into on of them

OpenStudy (kenljw):

Example 2x^2 + 4x +2 = 2(x^2 + 2x + 1) = 2(x+1)(x+1)=(2x+2)(x+1)

OpenStudy (kenljw):

If ask to solve the quadratic, a well divide out of equation and you get (x+1)^2